OpenStudy (anonymous):

which of the following choices is the standard deviation of sample shown here? 17,18,19,20,21 A,19 B,2 C,2.2 D, ^2 E,^2.5

OpenStudy (anonymous):

It is sample sd and so, division is by 4. The answer is D

OpenStudy (mathmate):

Do you mean \[D, \sqrt 2\ \ E, \sqrt{2.5})\]

OpenStudy (anonymous):

yea im pretty sure thats what they meant. its the only thing that makes sense.

OpenStudy (anonymous):

YEA HOW DO YOU DO THAT ???

OpenStudy (mathmate):

BTW, division by 4 does not give D.

OpenStudy (anonymous):

IT WOULD BE E^2.5

OpenStudy (mathmate):

You would type \ [\ s q r t 2\ ] but skip the spaces after the first and last \

OpenStudy (anonymous):

i meant E. sorry

OpenStudy (mathmate):

or better, go to the latex practising forum at openstudy.

OpenStudy (mathmate):

\[ \sqrt{\frac{1^2+2^2+2^2+1^2}{4}}=\sqrt{5/2}=\sqrt{2.5}\]

OpenStudy (anonymous):

THANKS

OpenStudy (mathmate):

You're welcome! :)